OpenStudy (anonymous):

If the emf produced in a wire is 0.88 volts and the wire moves perpendicular to a magnetic field of strength 0.075 newtons/amperes·meter at a speed of 4.20 meters/second, what is the length of the wire in the magnetic field?

OpenStudy (anonymous):

emf = -d(flux)/dt flux = B A, field times area normal to field B is constant but A is changing, as L sweeps along, dA/dt = L (4.20 m/s) emf = - B dA/dt